OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.

OpenStudy (anonymous):

f(x) = −one eighth (x − 2)^2 + 6 f(x) = one eighth (x − 2)^2 + 6 f(x) = −one eighth (x + 2)^2 + 8 f(x) = one eighth (x + 2)^2 + 8

OpenStudy (anonymous):

first example here will explain http://www.purplemath.com/modules/parabola3.htm

OpenStudy (amistre64):